# Uncertainty Estimation and Quantification for LLMs: A Simple Supervised Approach

### v1

@@ Abstract
Large language models (LLMs) are highly capable of many tasks but they can sometimes generate unreliable or inaccurate outputs. To tackle this issue, this paper studies the problem of uncertainty estimation and calibration for LLMs. We begin by formulating the uncertainty estimation problem for LLMs and then propose a supervised approach that takes advantage of the labeled datasets and estimates the uncertainty of the LLMs’ responses. Based on the formulation, we illustrate the difference between the uncertainty estimation for LLMs and that for standard ML models and explain why the hidden activations of the LLMs contain uncertainty information. Our designed approach effectively demonstrates the benefits of utilizing hidden activations for enhanced uncertainty estimation across various tasks and shows robust transferability in out-of-distribution settings. Moreover, we distinguish the uncertainty estimation task from the uncertainty calibration task and show that a better uncertainty estimation mode leads to a better calibration performance. In practice, our method is easy to implement and is adaptable to different levels of model transparency including black box, grey box, and white box, each demonstrating strong performance based on the accessibility of the LLM’s internal mechanisms.
**footnotetext: Equal contribution.footnotetext: Corresponding to: xiaocheng.li@imperial.ac.uk, guanting@unc.edu.
@@ 1 Introduction
Large language models (LLMs) have marked a significant milestone in the advancement of natural language processing (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022; Bubeck et al., 2023), showcasing remarkable capabilities in understanding and generating human-like text. However, one pressing issue for the LLMs is their propensity to hallucinate (Rawte et al., 2023) and generate misleading or entirely fabricated information that can significantly undermine their trustworthiness and reliability. The task of uncertainty estimation has then emerged to be an important problem, where an uncertainty estimation model can be used to determine the confidence levels of LLMs’ outputs. While the problem of uncertainty estimation and calibration has seen considerable development within the general machine learning and deep learning domains (Abdar et al., 2021; Gawlikowski et al., 2023), we see less development in the domain of LLMs. One of the major challenges is the difference in the format of the output: while machine learning and deep learning typically involve fixed-dimensional outputs, natural language generation (NLG) tasks central to LLM applications require handling variable outputs that carry semantic meanings, and it is unclear whether the uncertainty estimation for NLG should target for (fixed-dimensional) token level or (variable-dimensional) sentence/semantic-level uncertainty.
Existing uncertainty estimation approaches for LLMs usually involve designing uncertainty metrics for their outputs. For black-box LLMs, these metrics are computed by examining aspects like the generated outputs’ consistency, similarity, entropy, and other relevant characteristics (Lin et al., 2023; Manakul et al., 2023; Kuhn et al., 2023). Given the complexity of LLMs’ underlying architectures, semantic information may be diluted when processing through self-attention mechanisms and during token encoding/decoding. To address this issue, a growing stream of literature argues that hidden layers’ activation values within the LLMs offer insights into the LLMs’ knowledge and confidence (Slobodkin et al., 2023; Ahdritz et al., 2024; Duan et al., 2024). Based on this argument, white-box LLMs, which allow access to more of LLMs’ inner values, such as logits and hidden layers, are believed to have the capacity to offer a more nuanced understanding and improved uncertainty estimation results (Verma et al., 2023; Chen et al., 2024; Plaut et al., 2024).
@@
Our study embarks on a systematic evaluation of the integration of LLMs’ hidden states within the uncertainty estimation framework in a supervised manner. We aim to assess the tangible benefits this approach may offer in improving the confidence, reliability, transferability, and practicality of LLMs’ uncertainty estimation. By examining the impact of incorporating the knowledge of these hidden states into the uncertainty estimation methods, our results seek to clarify whether and how the internal states of LLMs can contribute to uncertainty estimation, paving the way for more trustworthy and dependable LLMs.
@@
  • We formulate the problem of uncertainty estimation for LLMs. Building on this formulation, we explore the nuanced distinctions between uncertainty estimation and calibration for LLMs, and theoretically demonstrate how uncertainty estimation for LLMs differs from that in traditional ML models. Moreover, our theoretical analysis suggests that the existing architecture and training procedures of LLMs require additional features from their input and output for improved uncertainty estimation.
  • Motivated by the above findings, we propose a supervised method for the uncertainty estimation problem. Specifically, the method aims to train an uncertainty estimation function that maps the hidden activations of the LLMs, and the probability-related information of the LLMs (when generating the response) to an uncertainty score that captures the LLM’s confidence about its response. This supervised approach is systematically designed for straightforward implementation and broad applicability, suitable for black-box, grey-box, and white-box LLMs.
  • We conduct numerical experiments on NLP tasks including question answering, machine translation, and multiple choice to evaluate the performance of our proposed approach against existing benchmarks. When designing these experiments, we measure the performance of our approach across LLMs with different accessibility mode (black-box, grey-box, and white-box), and evaluate them in both in-distribution and out-of-distribution test datasets. The results demonstrate that by leveraging hidden activations of LLMs, our method consistently extracts additional knowledge from these models to enhance uncertainty estimation across various NLP tasks. These findings provide insights into the working mechanism of the uncertainty estimation method and its robustness and transferability.
@@ 2 Problem Setup
Here the probability distribution $p_{\theta}$ denotes the distribution (over vocabulary $\mathcal{V}$) as the LLM’s output, and $\theta$ encapsulates all the parameters of the LLM. The conditional part includes the prompt $\bm{x}$ and all the tokens $y_{1},y_{2},...,y_{j-1}$ generated preceding the current position. We allow both the prompt and the response to have a variable length of $k$ and $m.$
We consider using the LLM for some downstream NLP tasks such as question answering, multiple choice, and machine translation. Such a task usually comes with an evaluation/scoring function that evaluates the quality of the generated response
$s(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1].$
For each pair of $(\bm{x},\bm{y}),$ the evaluation function rates the response $\bm{y}$ with the score $z\coloneqq s(\bm{y}_{\text{true}},\bm{y})$ where $\bm{y}_{\text{true}}$ is the true response for the prompt $\bm{x}$. The true response $\bm{y}_{\text{true}}$ is usually decided by factual truth, humans, or domain experts. It does not hurt to assume a larger score represents a better answer; $z=1$ indicates a perfect answer, while $z=0$ says the response $\bm{y}$ is off the target.
@@
where the expectation on the right-hand side is taken with respect to the (possible) randomness of the true response $\bm{y}_{\text{true}}$. We emphasize two points on this task definition:
  • The uncertainty function $g$ takes the prompt $\bm{x}$ and $\bm{y}$ as its inputs, which has the following implications. First, the true and predicted uncertainty score can and should depend on the specific realization of the response $\bm{y}$. Second, the uncertainty function $g$ does not require the true response $\bm{y}_{\text{true}}$ as the input. Since true response data are often limited and typically only available from labeled datasets or human experts, this design broadens the applicability of the function across various settings, enhancing its practical usage.
  • The uncertainty score function $g$, in the language of uncertainty calibration of ML models (Guo et al., 2017; Abdar et al., 2021), is defined in an individual (conditional) sense. That is, the predicted score $g(\bm{x},\bm{y})$ is hopefully precise and matches the true score on an individual level (for each prompt-response pair) but not on the population level (for a distribution of prompt-response pairs). With such a definition of the uncertainty quantification task, the score can be useful in that a well-calibrated score guides the extent to which the users should trust the response.
In the following sections, we will first describe our method as a generic framework of uncertainty estimation that formulates the problem as a supervised task and utilizes the available labeled dataset. Then we discuss how this approach can be effectively integrated in the state-of-the-art LLMs, and present the empirical experiments and findings. Finally, we discuss the relationship between our method and the existing literature on fine-tuning LLM and hallucination detection.
@@ 3 Uncertainty Estimation via Supervised Calibration
In Section 3.1, we present our method of supervised calibration as a post hoc procedure to estimate the uncertainty of the LLM’s responses.
@@ 3.1 Supervised calibration
$\mathcal{D}_{\text{raw}}$ can be generated based on a labeled dataset for the tasks we consider. Here $\bm{x}_{i}=(x_{i,1},...,x_{i,k_{i}})$ and $\bm{y}_{i}=(y_{i,1},...,y_{i,m_{i}})$ denote the prompt and the corresponding LLM’s response, respectively. $\bm{y}_{i,\text{true}}$ denotes the true response (that comes from the labeled dataset) of $\bm{x}_{i}$, and $s(\bm{y}_{i},\bm{y}_{i,\text{true}})$ assigns a score for the response $\bm{y}_{i}$ based on the true answer $\bm{y}_{i,\text{true}}$. We remark that when generating the dataset $\mathcal{D}_{\text{raw}}$ with the LLM, there can be multiple samples with the same prompt for tasks such as question answering and machine translation. That is, we can generate multiple samples with $\bm{x}_{i}\equiv\bm{x}$, but these samples may have different responses $\bm{y}_{i}$’s due to the probabilistic nature of the LLM. Meanwhile, because of the same $\bm{x}_{i}$, these samples have the same true response $\bm{y}_{i,\text{true}}$. Each different realization of $\bm{y}_{i}$ constitutes one meaningful sample for the uncertainty estimation task.
The next step is to formulate a supervised learning task based on the dataset $\mathcal{D}_{\text{raw}}$. Specifically, we construct the following dataset
@@
where $z_{i}\coloneqq s(\bm{y}_{i},\bm{y}_{i,\text{true}})\in[0,1]$ denotes the target score to be predicted. The vector $\bm{v}_{i}$ summarizes useful features for the $i$-th sample based on $(\bm{x}_{i},\bm{y}_{i})$ for the prediction task. In this light, a supervised learning task on the dataset $\mathcal{D}_{\text{un}}$ corresponds exactly to the definition of the uncertainty estimation task in (1).
Now we discuss how we construct the feature vector $\bm{v}_{i}$ based on $(\bm{x}_{i},\bm{y}_{i})$. We mainly consider features from two sources.
  • White-box features: LLM’s hidden-layer activations. We feed $(\bm{x}_{i},\bm{y}_{i})$ as input into a LLM, and extract the corresponding hidden layers’ activations of the LLM. Given the architecture of the decoder-based transformer, the consensus is that the activations of the last token (of an input sequence) should contain the most information (Azaria and Mitchell, 2023; Chen et al., 2024). Thus we use the hidden layers’ activation of the last token of $(\bm{x}_{i},\bm{y}_{i})$ as part of $\bm{v}_{i}$. We defer more discussions on the choice of the layer to Section 4.3.
  • Grey-box features: Entropy- or probability-related outputs. The entropy of a discrete distribution $p$ over the vocabulary $\mathcal{V}$ is defined by
    $H(p)\coloneqq-\sum_{v\in\mathcal{V}}p(v)\log\left(p(v)\right).$
    For a prompt-response pair $(\bm{x},\bm{y})=(x_{1},...,x_{k},y_{1},...,y_{m})$, we consider as the features the entropy at each token such as $H(p_{\theta}(\cdot|x_{1},...,x_{j-1}))$ and $H(p_{\theta}(\cdot|\bm{x},y_{1},...,y_{j-1}))$ where $p_{\theta}$ denotes the LLM.
We defer more discussions on feature construction to Appendix A.1. Also, we note that another direct feature for predicting $z_{i}$ is to ask the LLM “how certain it is about the response” and incorporate its response to this question as a feature for predicting $z_{i}$ (Tian et al., 2023). This is naturally a viable option, and there are also other sources of features that can be incorporated into $\bm{v}_{i}$ such as the hidden activations for the second last token. We do not try to exhaust all the possible features, and the aim of our paper is more about formulating the uncertainty estimation for the LLMs as a supervised task and understanding how the internal states of the LLM encode uncertainty. To the best of our knowledge, our paper is the first one to do so. Specifically, the above formulation aims for the following two outcomes:
  • We train an uncertainty model $\hat{g}(\bm{v}_{i})$ that predicts $z_{i}.$ Based on the dataset $\mathcal{D}_{\text{un}}=\left\{(\bm{v}_{i},z_{i})\right\}_{i=1}^{n}$, one can learn a model $\hat{g}$ that predicts the uncertainty $z_{i}$ with summarized features $\bm{v}_{i}$. Intuitively, this approach will give better performance (because of its supervised nature and the reason presented in Section 3.3) than the ad-hoc/black-box methods such as the entropy-based method (Malinin and Gales, 2021; Kuhn et al., 2023), similarity-based method (Lin et al., 2022), etc., which is indeed verified by the numerical experiments in the next section. Thus this result underscores the usefulness of the supervised dataset for predicting uncertainty which is more aligned with the canonical uncertainty quantification methods.
  • We aim to answer the question of whether the hidden layers carry the uncertainty information. We note that using hidden layer information is not a mainstream approach for the existing uncertainty estimation literature on ML models. This highlights the difference between the uncertainty estimation for classic deep learning models and that for LLMs. We provide some theoretical insights in the following sections and show by theoretical insights and empirical experiments why the supervised approach can extract additional knowledge of LLMs to enhance uncertainty estimation.
@@
So far in this paper, we focus on the uncertainty estimation task which aims to predict whether the LLM makes mistakes in its response or not. There is a different but related task known as the uncertainty calibration problem. In comparison, the uncertainty calibration aims to ensure that the output from the uncertainty estimation model such as $\hat{g}$ in above conveys a probabilistic meaning. In this sense, our paper is more focused on addressing the predictability of the LLM’s uncertainty, and we find that our supervised formulation and the LLM’s hidden activations are helpful in such a prediction task. In terms of the uncertainty calibration aspect, our developed uncertainty estimation model is compatible with all the recalibration methods for ML models in the literature of uncertainty calibration. And intuitively, a better uncertainty estimation/prediction will lead to a better-calibrated uncertainty model, which is also verified in our numerical experiments in Section 4.4.
@@ 3.3 Why hidden layers as features?
[…]
$f^{*}=\operatorname*{arg\,max}_{f\in\mathcal{F}}\mathbb{E}\left[l(Y,f(\bm{X}))\right]$
@@
The proposition is not technical and it can be easily proved by using the structure of $f^{*}(\bm{X})$. It states a nice (probably well-known) property of the cross-entropy loss. Specifically, the function learned under the cross-entropy loss coincides with the Bayes optimal classifier. Note that this is contingent on two requirements. First, the function class $\mathcal{F}$ is the measurable function class. Essentially, it suffices to require the function class $\mathcal{F}$ to be rich enough to cover the Bayesian optimal classifier, also known as the realizability condition. With the capacity of the large vision and NLP models, we can think this requirement to be (at least approximately) met. Second, it requires the function $f^{*}$ learned through the population loss rather than the empirical loss/risk consisting of the training samples. With a large size of training data, we can think this requirement also to be (approximately) met.
The proposition also states one step further on conditional independence $Y\perp\bm{X}\ |\ f^{*}(\bm{X})$. This means all the information related to the label $Y$ that is contained in $\bm{X}$ is summarized in the prediction function $f^{*}.$ While the numeric situation cannot be captured by such a simple theory (due to these two requirements), it provides insights into why the existing uncertainty quantification/calibration/recalibration methods do not utilize much of the original feature $\bm{X}$ or the hidden-layer activations. Specifically, when a prediction model $\hat{f}:\mathbb{R}^{d}\rightarrow[0,1]$ is well-trained, the predicted score $\hat{f}(\bm{X})$ should capture all the information about the true label $Y$ contained in the features $\bm{X}$, and there is no need to get the features $\bm{X}$ re-involved in the recalibration procedure when one adjusts the prediction model $\hat{f}$ to meet some calibration objectives. This indeed explains why the classic uncertainty quantification and calibration methods only work with the predicted score $\hat{f}(\bm{X})$ for re-calibration, including Platt scaling (Platt et al., 1999), isotonic regression (Zadrozny and Elkan, 2002), temperature scaling (Guo et al., 2017), etc.
@@
$\tilde{f}=\operatorname*{arg\,max}_{f\in\mathcal{F}}\mathbb{E}\left[\tilde{l}(Y,f(\bm{X}))\right],$
@@
Proposition 3.1 and Corollary 3.2 together illustrate the difference between uncertainty estimation for a traditional ML model and that for LLMs. For the traditional ML models, the cross-entropy loss which is commonly used for training the model is aligned toward the uncertainty calibration objective. When it comes to uncertainty estimation for LLMs, the label $Y$ can be viewed as the binary variable for whether the LLM’s response is correct, and the features $\bm{X}$ represent some features extracted from the prompt-response pair. Meanwhile, the LLMs are often pre-trained with some other loss functions (for example, the negative log-likelihood loss for next-token prediction), and this causes a misalignment between the model pre-training and the uncertainty quantification task. The consequence is the collapse of conditional independence in Corollary 3.2. The collapse is of a deeper extent than the violation of the two requirements as discussed above. Consequently, the original features, the prompt-response pair, may and should (in theory) contain information about the uncertainty score $Y$ that cannot be fully captured by $\tilde{f}(\bm{X})$. This justifies why we formulate the uncertainty estimation task as the previous subsection and take the hidden-layer activations as features to predict the uncertainty score; it also explains why we do not see much similar treatment in the mainstream uncertainty quantification literature.

### v2

@@ Abstract
In this paper, we study the problem of uncertainty estimation and calibration for LLMs. We first formulate the uncertainty estimation problem for LLMs and then propose a supervised approach that takes advantage of the labeled datasets and estimates the uncertainty of the LLMs’ responses. Based on the formulation, we illustrate the difference between the uncertainty estimation for LLMs and that for standard ML models and explain why the hidden neurons of the LLMs may contain uncertainty information. Our designed approach demonstrates the benefits of utilizing hidden activations to enhance uncertainty estimation across various tasks and shows robust transferability in out-of-distribution settings. We distinguish the uncertainty estimation task from the uncertainty calibration task and show that a better uncertainty estimation mode leads to a better calibration performance. Furthermore, our method is easy to implement and adaptable to different levels of model accessibility including black box, grey box, and white box.
**footnotetext: Equal contribution.
@@ 1 Introduction
Large language models (LLMs) have marked a significant milestone in the advancement of natural language processing (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022; Bubeck et al., 2023), showcasing remarkable capabilities in understanding and generating human-like text. However, one pressing issue for the LLMs is their propensity to hallucinate (Rawte et al., 2023) and generate misleading or entirely fabricated information that can significantly undermine their trustworthiness and reliability. The task of uncertainty estimation has then emerged to be an important problem that aims to determine the confidence levels of LLMs’ outputs. While the problem of uncertainty estimation and calibration has seen considerable development within the general machine learning and deep learning domains (Abdar et al., 2021; Gawlikowski et al., 2023), we see less development in the domain of LLMs. One of the major challenges is the difference in the format of the output: while machine learning and deep learning typically involve fixed-dimensional outputs, natural language generation (NLG) tasks central to LLM applications require handling variable outputs that carry semantic meanings.
Existing uncertainty estimation approaches for LLMs usually involve designing uncertainty metrics for their outputs. For black-box LLMs, these metrics are computed by examining aspects like the generated outputs’ consistency, similarity, entropy, and other relevant characteristics (Lin et al., 2023; Manakul et al., 2023; Kuhn et al., 2023; Hou et al., 2023). Given the complexity of LLMs’ underlying architectures, semantic information may be diluted when processing through self-attention mechanisms and during token encoding/decoding. To address this issue, a growing stream of literature argues that hidden layers’ activation values within the LLMs offer insights into the LLMs’ knowledge and confidence (Slobodkin et al., 2023; Ahdritz et al., 2024; Duan et al., 2024). Based on this argument, white-box LLMs, which allow access to more of LLMs’ inner values, such as logits and hidden layers, are believed to have the capacity to offer a more nuanced understanding and improved uncertainty estimation results (Verma et al., 2023; Chen et al., 2024; Plaut et al., 2024).
@@
- First, we formulate the problem of uncertainty estimation for LLMs. We explore the nuanced distinctions between uncertainty estimation and calibration for LLMs, and theoretically demonstrate how uncertainty estimation for LLMs differs from that in traditional ML models.
- Second, motivated by the above findings, we propose a supervised method for the uncertainty estimation problem. Specifically, the method aims to train an uncertainty estimation function that maps the hidden activations of the LLMs, and the probability-related information of the LLMs (when generating the response) to an uncertainty score that captures the LLM’s confidence about its response. This supervised approach is systematically designed for straightforward implementation and broad applicability, suitable for black-box, grey-box, and white-box LLMs.
- Third, we conduct numerical experiments on NLP tasks including question answering, machine translation, and multiple choice to evaluate our proposed approach against existing benchmarks. We systematically investigate the working mechanism of uncertainty quantification.
@@ 2 Problem Setup
Here the probability distribution $p_{\theta}$ denotes the distribution (over vocabulary $\mathcal{V}$) as the LLM’s output, and $\theta$ encapsulates all the parameters of the LLM. The conditional part includes the prompt $\bm{x}$ and all the tokens $y_{1},y_{2},...,y_{j-1}$ generated preceding the current position.
We consider using the LLM for some downstream NLP tasks such as question answering, multiple choice, and machine translation. Such a task usually comes with an evaluation/scoring function that evaluates the quality of the generated response $s(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1].$ For each pair of $(\bm{x},\bm{y}),$ the evaluation function rates the response $\bm{y}$ with the score $z\coloneqq s(\bm{y}_{\text{true}},\bm{y})$ where $\bm{y}_{\text{true}}$ is the true response for the prompt $\bm{x}$. The true response $\bm{y}_{\text{true}}$ is usually decided by factual truth, humans, or domain experts. It does not hurt to assume a larger score represents a better answer; $z=1$ indicates a perfect answer, while $z=0$ says the response $\bm{y}$ is off the target.
@@
where the expectation on the right-hand side is taken with respect to the (possible) randomness of the true response $\bm{y}_{\text{true}}$. We emphasize two points on this task definition: The uncertainty function $g$ takes the prompt $\bm{x}$ and $\bm{y}$ as its inputs. This implies (i) the true and predicted uncertainty score can and should depend on the specific realization of the response $\bm{y}$ and (ii) the uncertainty function $g$ does not require the true response $\bm{y}_{\text{true}}$ as the input.
The uncertainty score function $g$, in the language of uncertainty calibration of ML models (Guo et al., 2017; Abdar et al., 2021), is defined in an individual (conditional) sense. That is, the predicted score $g(\bm{x},\bm{y})$ is hopefully precise and matches the true score on an individual level (for each prompt-response pair) but not on the population level (for a distribution of prompt-response pairs).
@@ 3.1 Supervised calibration
$\mathcal{D}_{\text{raw}}$ can be generated based on a labeled dataset for the tasks we consider. Here $\bm{x}_{i}=(x_{i,1},...,x_{i,k_{i}})$ and $\bm{y}_{i}=(y_{i,1},...,y_{i,m_{i}})$ denote the prompt and the corresponding LLM’s response, respectively. $\bm{y}_{i,\text{true}}$ denotes the true response (that comes from the labeled dataset) of $\bm{x}_{i}$, and $s(\bm{y}_{i},\bm{y}_{i,\text{true}})$ assigns a score for the response $\bm{y}_{i}$ based on the true answer $\bm{y}_{i,\text{true}}$.
The next is to formulate a supervised learning task based on $\mathcal{D}_{\text{raw}}$. Specifically, we construct
@@
where $z_{i}\coloneqq s(\bm{y}_{i},\bm{y}_{i,\text{true}})\in[0,1]$ denotes the target score to be predicted. The vector $\bm{v}_{i}$ summarizes useful features for the $i$-th sample based on $(\bm{x}_{i},\bm{y}_{i})$. That is, a supervised learning task on the dataset $\mathcal{D}_{\text{un}}$ coincides exactly to the definition of the uncertainty estimation task in (1). We mainly consider features from two sources when constructing $\bm{v}_{i}$ based on $(\bm{x}_{i},\bm{y}_{i})$:
White-box features: LLM’s hidden-layer activations. We feed $(\bm{x}_{i},\bm{y}_{i})$ as input into a LLM, and extract the corresponding hidden layers’ activations of the LLM.
Grey-box features: Entropy- or probability-related outputs. The entropy of a discrete distribution $p$ over the vocabulary $\mathcal{V}$ is defined by $H(p)\coloneqq-\sum_{v\in\mathcal{V}}p(v)\log\left(p(v)\right).$ For a prompt-response pair $(\bm{x},\bm{y})=(x_{1},...,x_{k},y_{1},...,y_{m})$, we consider as the features the entropy at each token such as $H(p_{\theta}(\cdot|x_{1},...,x_{j-1}))$ and $H(p_{\theta}(\cdot|\bm{x},y_{1},...,y_{j-1}))$ where $p_{\theta}$ denotes the LLM. We defer more discussions on feature construction to Appendix C. There can be other useful features such as to ask the LLM “how certain it is about the response” (Tian et al., 2023). We do not try to exhaust all the possiblility, and the aim of our paper is more about formulating the uncertainty estimation for the LLMs as a supervised task and understanding how the internal states of the LLM encode uncertainty. To the best of our knowledge, our paper is the first one to do so. Specifically, the above formulation aims for the following two outcomes: (i) an uncertainty model $\hat{g}(\bm{v}_{i})$ that predicts $z_{i}$ and (ii) knowing whether the hidden layers carry the uncertainty information.
@@
So far in this paper, we focus on the uncertainty estimation task which aims to predict whether the LLM makes mistakes in its response or not. There is a different but related task known as the uncertainty calibration problem. In comparison, the uncertainty calibration aims to ensure that the output from the uncertainty estimation model such as $\hat{g}$ in above conveys a probabilistic meaning. In this sense, our paper is more focused on addressing the predictability of the LLM’s uncertainty, and we find that our supervised formulation and the LLM’s hidden activations are helpful in such a prediction task. In terms of the uncertainty calibration aspect, our developed uncertainty estimation model is compatible with all the recalibration methods for ML models in the literature of uncertainty calibration. And intuitively, a better uncertainty estimation/prediction will lead to a better-calibrated uncertainty model, which is also verified in our numerical experiments in Appendix B.
@@ 3.3 Why hidden layers as features?
Let $\mathcal{F}$ be the class of measurable function that maps from $\mathbb{R}^{d}$ to $[0,1]$. Under the cross-entropy loss $l(y,\hat{y})=y\log(\hat{y})+(1-y)\log(1-\hat{y})$, the function $f^{*}$ that minimizes the loss
$f^{*}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\mathbb{E}\left[l(Y,f(\bm{X}))\right]$
@@
The proposition is not technical and it can be easily proved by using the structure of $f^{*}(\bm{X})$. It states a nice (probably well-known) property of the cross-entropy loss. Specifically, the function learned under the cross-entropy loss coincides with the Bayes optimal classifier. Note that this is contingent on two requirements. First, the function class $\mathcal{F}$ is the measurable function class. Second, it requires the function $f^{*}$ learned through the population loss rather than the empirical loss/risk consisting of the training samples. The proposition also states one step further on conditional independence $Y\perp\bm{X}\ |\ f^{*}(\bm{X})$. This means all the information related to the label $Y$ that is contained in $\bm{X}$ is summarized in the prediction function $f^{*}.$ While the numeric situation cannot be captured by such a simple theory (due to these two requirements), it provides insights into why the existing uncertainty quantification/calibration/recalibration methods do not utilize much of the original feature $\bm{X}$ or the hidden-layer activations. Specifically, when a prediction model $\hat{f}:\mathbb{R}^{d}\rightarrow[0,1]$ is well-trained, the predicted score $\hat{f}(\bm{X})$ should capture all the information about the true label $Y$ contained in the features $\bm{X}$, and there is no need to get the features $\bm{X}$ re-involved in the recalibration procedure. This indeed explains why the classic uncertainty quantification and calibration methods only work with the predicted score $\hat{f}(\bm{X})$ for re-calibration, including Platt scaling (Platt et al., 1999), isotonic regression (Zadrozny and Elkan, 2002), temperature scaling (Guo et al., 2017), etc.
@@
$\tilde{f}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\mathbb{E}\left[\tilde{l}(Y,f(\bm{X}))\right],$
@@
Proposition 3.1 and Corollary 3.2 together illustrate the difference between uncertainty estimation for a traditional ML model and that for LLMs. For the traditional ML models, the cross-entropy loss which is commonly used for training the model is aligned toward the uncertainty calibration objective. When it comes to uncertainty estimation for LLMs, the label $Y$ can be viewed as the binary variable for whether the LLM’s response is correct, and the features $\bm{X}$ represent some features extracted from the prompt-response pair. Meanwhile, the LLMs are often pretrained with some other loss functions (for example, the negative log-likelihood loss for next-token prediction), and this causes a misalignment between the model pre-training and the uncertainty quantification task. Consequently, the original features, the prompt-response pair, may and should (in theory) contain information about the uncertainty score $Y$ that cannot be fully captured by $\tilde{f}(\bm{X})$. This justifies why we formulate the uncertainty estimation task as the previous subsection and take the hidden-layer activations as features to predict the uncertainty score; it also explains why we do not see much similar treatment in the mainstream uncertainty quantification literature.